if the tempeture of 50.0L of a gas at 40.0 degrees C falls by 10.0 degrees C what is the new volume of the gas if the pressure is constant?
1
Expert's answer
2013-05-08T09:25:33-0400
For an ideal gas at constant pressure, the volume is directly proportional to its temperature.
V1/T1 = V2/T2
where
V is volume
T is temperature (K)
V1 = 50.0 L
T1 = 313 K
If the temperature falls by 10.0 C, so T2 will be = 40 - 10 = 30 C or 303 K
Therefore V2 will be
V2 = V1*T2/T1
V2 = 50.0 L * 303 K / 313 K = 48.40 L
Answer: V2 = 48.40 L